OpenStudy (anonymous):

A particle moves along the curve f(x). As it reaches the point (3; 2), the y-coordinate is increasing by 1.2 cm/sec. At what rate is the x-coordinate changing at that instant?

OpenStudy (anonymous):

y=\[\sqrt[3]{x^{2}-1}\]

OpenStudy (anonymous):

here they actually give you the formula

OpenStudy (anonymous):

I take the derivative.

OpenStudy (anonymous):

but then what?

OpenStudy (anonymous):

yes, don't forget that \(x\) is a function of \(t\) also, your answer should have a \(x'\) in it

OpenStudy (anonymous):

oh, so I just plug in 1.2 to x' and 3 into x and get y'?

OpenStudy (anonymous):

at the end, yes you should get as a first step \[y'=\frac{2xx'}{3\sqrt{x^2-1}^2}\]

OpenStudy (anonymous):

I got 0.6